# Evolutionary Algorithm with Detection Region Method for Constrained Multi-Objective Problems with Binary Constraints

### v1

@@ 5.2 Ablation Study Results Analysis
5.3 Experimental Results on real-world applications
@@
where $F_{1}$ represents the total mileage of the vehicle, $N_{c}$ is the number of customers, $c_{ij}$ is the distance between the $i$th customer and the $j$th customer; when $i=0$, $c_{0j}$ represents the distance from the warehouse to the $j$th customer, when $j=0$, $c_{i0}$ represents the distance from the $i$th customer to the warehouse, and $x_{ij}$ represents whether there is a path between the $i$th customer and the $j$th customer, where $x_{ij}$ = 1 means there is a path, otherwise $x_{ij}$ = 0. $F_{2}$ is the load variance between vehicles, where $M_{v}$ is the number of vehicles used, $d_{m}$ represents the load of the m-th vehicle, and $\overline{d}$ is the average load of $M_{v}$ vehicles. In addition, the MO-CVRP needs to satisfy seven rule-based constraints. The specific equation definitions can be found in the original paper [56]. In this paper, all rule-based constraints in MO-CVRP can only return satisfied or unsatisfied. The problem scenario dataset we use is the widely used CVRP dataset ”Classic M” [57] [58]. This dataset contains 5 problems. Taking ”M-n101-k10” as an example, ”n101” represents 101 customers, and ”k10” represents 10 vehicles. The number of evaluations for all problems is set to 10000 times, and the population size is set to 100. Table 3 shows the HV results of our DRMCMO and other comparison algorithms on this real problem. It can be seen that DRMCMO achieves the best results on all problems and significantly outperforms other algorithms. In the comparison of HV convergence curves shown in Fig. 10, DRMCMO exhibits the fastest convergence speed. Additionally, the Friedman ranking in Fig. 8 shows that DRMCMO surpasses other leading CMOEAs by a substantial margin. The experimental results fully demonstrate the competitiveness of our proposed DRMCMO in handling real-world CMOP/BC.
6 Conclusion and Future Work
@@ Acknowledgment
This work was supported by the Scientific Research Project of Xiang Jiang Lab (22XJ02003), the University Fundamental Research Fund (23-ZZCX-JDZ-28), the National Science Fund for Outstanding Young Scholars (62122093), the National Natural Science Foundation of China (72071205), the science and technology innovation Program of Hunan Province (ZC23112101-10), and the Hunan Natural Science Foundation Regional Joint Project (2023JJ50490).
The authors would like to thank the support by the COSTA: complex system optimization team of the College of System Engineering at NUDT.

### v2

@@ V-B Ablation Study Results Analysis
V-C Experimental Results on real-world applications
@@
where $F_{1}$ represents the total mileage of the vehicle, $N_{c}$ is the number of customers, $c_{ij}$ is the distance between the $i$th customer and the $j$th customer; when $i=0$, $c_{0j}$ represents the distance from the warehouse to the $j$th customer, when $j=0$, $c_{i0}$ represents the distance from the $i$th customer to the warehouse, and $x_{ij}$ represents whether there is a path between the $i$th customer and the $j$th customer, where $x_{ij}$ = 1 means there is a path, otherwise $x_{ij}$ = 0. $F_{2}$ is the load variance between vehicles, where $M_{v}$ is the number of vehicles used, $d_{m}$ represents the load of the m-th vehicle, and $\overline{d}$ is the average load of $M_{v}$ vehicles. In addition, the MO-CVRP needs to satisfy seven rule-based constraints. The specific equation definitions can be found in the original paper [56]. In this paper, all rule-based constraints in MO-CVRP can only return satisfied or unsatisfied. The problem scenario dataset we use is the widely used CVRP dataset ”Classic M” [57] [58]. This dataset contains 5 problems. Taking ”M-n101-k10” as an example, ”n101” represents 101 customers, and ”k10” represents 10 vehicles. The number of evaluations for all problems is set to 10000 times, and the population size is set to 100. Table III shows the HV results of our DRMCMO and other comparison algorithms on this real problem. It can be seen that DRMCMO achieves the best results on all problems and significantly outperforms other algorithms. In the comparison of HV convergence curves shown in Fig. 10, DRMCMO exhibits the fastest convergence speed. Additionally, the Friedman ranking in Fig. 8 shows that DRMCMO surpasses other leading CMOEAs by a substantial margin. The experimental results fully demonstrate the competitiveness of our proposed DRMCMO in handling real-world CMOP/BC.
VI Conclusion and Future Work
@@ Acknowledgment
This work was supported by the Scientific Research Project of Xiang Jiang Lab (22XJ02003), the University Fundamental Research Fund (23-ZZCX-JDZ-28), the National Science Fund for Outstanding Young Scholars (62122093), the National Natural Science Foundation of China (72421002), the National University of Defense Technology Youth Innovation Science Fund Project (ZK25-62), the science and technology innovation Program of Hunan Province (ZC23112101-10), and the Hunan Natural Science Foundation Regional Joint Project (2023JJ50490). The authors would like to thank the support by the COSTA: complex system optimization team of the College of System Engineering at NUDT.